OpenStudy (anonymous):

The region between the graphs of y=x^2 and y=6x is rotated around the line x=6. The volume of the resulting solid is ?

OpenStudy (anonymous):

h(x) = f(x) - g(x) = x² - 6x Vol.Rotate = 2*Pi * Integral( h(x)² ) dx

OpenStudy (callisto):

1. Find their intersection points. => Put y=x^2 into y=6x => x(x-6)=0 => x=0 or x=6 2. \[ volume=2π \int_{0}^{6} (6-x)(6x) - (6-x)(x^2) dx\] Can you do it from here? (PS: I'm not sure if it is correct :S)

OpenStudy (anonymous):

Yup I got it from here, Thanks a lot!

OpenStudy (callisto):

Welcome :) Note: This time, I still use shell method to do it..